A new project is expected to have a FCF of $6M one year from today. The yearly cashflows will increase by 3% per year, forever. You’ve found a comparable firm that takes on similar projects and has an expected return on equity of 10%, an expected return on debt of 6%, and a D/V ratio of 0.3. You’ve decided to keep a constant D/E ratio of 1.0 for the project. Your cost of debt is 4% and the corporate tax rate is 40%. What is the NPV of the project?
Answer #1

Expected FCF, FCF1 = $ 6 M

Constant growth rate, g = 3%

Cost of equity, ke = 10%

Cost of debt, kd = 4%

t = 40%

D/V, Wd = 0.5

We = 0.5

Cost of capital, WACC = 4% * 0.5* 0.6 + 10% * 0.5 = 6.2%

Terminal value of cash flows = 6*(1+3%)/(6.2% - 3%) = $ 193.125

NPV of cash flows = 6/(1+6.2%) + 193.125/(1+ 6.2%) = $ 187.5 M